OpenStudy (anonymous):

Geometry help please! :) Okay, so, the question is Find the area of the circle. Leave your answer in terms of pi. The diameter of the circle is 4.1 m. A. 4.2025 pi B. 8.405 pi C. 16.81 pi D. 11.2 pi If i'm not mistaken, area is pi x radius squared. And the radius is half of the diameter, so it'd be 2.05. I end up with 6.437, and that isn't an option. What am I doing wrong here? haha.

jimthompson5910 (jim_thompson5910):

leave pi as it is (ie don't replace it with 22/7, or 3.14, etc) Area: A = pi*r^2 A = pi*2.05^2 ... plug in r = 2.05 A = pi*4.2025 A = 4.2025pi

OpenStudy (anonymous):

Thank you! :)

jimthompson5910 (jim_thompson5910):

you're welcome